Be sure to answer all parts At 50.00 degree C, wh
Answer #1

We know the relation for fraction of collisons

ƒ=e^(-Ea/RT)

From given data, activation energy Ea= 84.60 kj/mole = 84600 J/mole

Temperature T = 50 +273 = 323 K

R = 8.314 J/mole.K

Let us plugin all values in the equation, we will get

f =e^(-84600/(8.314*323))
=2.08 x10^-14